# Microsoft cloud-based digitization workflow with rich metadata acquisition for cultural heritage objects

### v1

@@ 4 Objectives and requirements
Our ultimate goal is to create a knowledge base that adheres to FAIR principles [24], that is linked to general knowledge bases—such as Wikidata [5], Geonames [^4] https://www.geonames.org/., and YAGO [22]—as well as to other cultural heritage bases—including the pan-European Europeana project [6] and bases describing the specific collections, e.g., the Kalliope catalog [^5] https://kalliope-verbund.info/en/. of approximately 600,000 records from 950 institutions located in German-speaking countries—and that will enable and facilitate a wide range of digital humanities research, including tasks done or supported by artificial intelligence tools.
As highlighted in the Introduction, the objective of the work reported in this paper was to develop tools for metadata acquisition—which as part of further work will be used to build a knowledge graph—by domain experts and to combine them with existing systems and tools (described in Sect. 3) into a unified digitization workflow for cultural heritage collections. From the very beginning, the Berlin Collection analyzed in the DiHeLib project (see Sect. 3) was chosen as the playground for the work. Therefore, our research concerns practical challenges of massive digitization (approximately 300,000 documents) of manuscripts of immense historical value for the purpose of a digital library.
@@
    Tight time frame for tool development – four months were allocated for this task in the DiHeLib project,
@@
    Lack of human and financial resources in the project to develop dedicated software,
@@ 5 Methods and tools
    A set of Python scripts performing more detailed validation of metadata entered into the TCM and SEC (see Sect. 5.2),
@@ 5.1 Designing TKM, SEC, and MAP
The Excel-based form for entering document metadata is the key DiHeLib data structure. First, it was assumed that each unit (see Sect. 3) would correspond to one file, in which one of the sheets would be a metric containing the description of a unit as a whole (see Fig. 4), and the other would be used to collect metadata of individual documents represented by subsequent rows (see Fig. 5). In turn, columns of the latter sheet would represent successive fields in the document description. Thus, The Knowledge Matrix emerged. It was decided that each document would be classified into one of nine categories, a decision that determines the assignment of a particular set of fields to it. The list is a slightly modified version of the scheme used for the Varnhagen Collection [18] and comprises the following categories:
@@
Some basic fields are ascribed to all document types, e.g., “what is it / document title”, while other are associated with particular categories, e.g., “sender” and “recipient” are assigned to correspondence (categories 2 and 7), “issuer” is assigned to official documents (categories 4 and 5), and “author” to creative works (category 3). The description fields were further divided into mandatory (as “document number”) and optional (as “date remarks”). Detailed content restrictions were imposed on some of the fields, e.g., “date” field could only accept one of the specified date formats. The most important from the validation point of view (see Sect. 5.2) is “document number” filled with a string in format $x.y$, where $x$ is a digit of encoding document category of 1 to 9, and $y$ is a sequential number of a document within a unit.
During subsequent iterations, the need for separate acquisition of person- and place-related metadata appeared. The problem arose from the fact that numerous individuals (or geographical names) were referred to in the documents under various different names, stemming from the different spellings of the same surnames in different languages or the fact that a person (or institution) might have assumed different names at different stages of their life. This need led to the creation of the Standard Entries Catalog. The SEC consists of four Excel files. Two of them are the parts of the actual catalog, presenting information about persons (see Fig. 6) and locations, respectively. Only the SEC coordinator is able to enter new data or modify existing entries, while other experts have read-only access. The other two files serve as collectors of proposals from domain experts: they can write suggestions of new entries or extending or modifying the existing entries in both main SEC files. The TKM records link to specific SEC entries by providing the SEC ID in a dedicated column. The SEC records are linked to external databases by providing URLs (e.g. Geonames URLs for geographic names).
@@
Finally, when the TKM file is finished, the validation script (see Sect. 5.2) extends it automatically with a MAP sheet designed to store bindings between unit’s card numbers and their scans. A single page may sometimes require more than just two scans (one for its recto, another for its verso side). There may be one or more paste-in slips attached to the page that require separate scanning. A telling example of this problem is one of the pages from Alexander von Humboldt’s Kosmos manuscript whose scanning produced as many as 14 different files (see pp. 62-75 in [11]). Therefore, to enable linking particular documents with their scans, a system of mapping card numbers to scan files—done manually by librarians in a separate MAP sheet—and mapping units with their documents to specific folder in a folder tree where the files are stored was introduced.
@@ 5.2 Data validation scripts
Data entered through interfaces (Excel files) must be evaluated to ensure the quality of the metadata is as high as possible. Two programs that perform data validation have been developed for this purpose: 1) on the fly, as they are entered by the researcher into the TKM sheet (via an Office Script), and 2) after the sheet has been filled out completely – in this case, a full analysis is possible, which is technically impossible to perform at the Excel online script level (via a set of Python scripts). In both cases, the syntax of the input data and its completeness are analyzed. Program 2 additionally performs limited semantic validation including, but not limited to: verification whether the Author’s life span is not greater than 110 years, whether the entered dates actually existed in the calendar, whether the entered SEC IDs indeed exist in the SEC and correspond to the specified person/place. Feedback is provided to users through appropriate messages in a dedicated column A and cell background colors (in Program 1) and in the form of a report in a PDF file (in Program 2).
Both programs were dedicated to data validation, however, during the course of work, their scope of operation was extended with additional functionalities resulting from the current needs of the Project:
@@
    Updating the TKM with new columns and fields, performed by Program 1 (e.g. adding new fields that were not provided for earlier; the script performs this change without the need for additional user intervention);
@@
    Validation of SEC entries by Program 2 (an independent report is generated, visible to the SEC coordinator).
@@
For Program 2, a dedicated virtual server on the JU infrastructure was set up and configured to perform a synchronization with Sharepoint once a day, run the validation scripts, and perform the synchronization again to propagate the validation reports.
6 Digitization workflow in practice
@@
Data entry and validation are only the initial steps in an overall metadata acquisition process for digitization integrated with BJ’s procedures. Working closely with JL, through a series of iterations, existing processes at JL were first identified, and then a new complete process was developed using Excel input interfaces, the Microsoft 365 cloud, and JL’s existing infrastructure. Importantly, the process was subjected to practical evaluation through two pilots and two workshops with international guests, which allowed for its improvement, and confirmation of its correctness and usability for JL.
@@
Designing, organizing, and implementing the workflow process, and consequently its implementation in practice, was a major organizational challenge due to three key factors. Firstly, it included the activities of three different groups of people, partly working in different places, with different levels of IT competences, using different IT tools on a daily basis and coming from different work and organizational cultures: domain experts – humanities researchers, computer scientists – researchers, academic teachers, doctoral students and practitioners, and library workers. Additionally, the last of the mentioned groups was also highly differentiated internally. They included librarians specializing in working with manuscripts, as well as a library system administrator, a person specializing in publishing objects in a digital library, a person operating a scanner and an IT specialist responsible for the local network and, to some extent, servers. Secondly, the workflow was to be fed with different resources at different times: on the one hand, original manuscripts that require special protection and security procedures, and on the other hand, two types of digital documents: Excel files filled with metadata and scans of the manuscripts. Thirdly, the Jagiellonian Library did not have a ready-made tool capable of comprehensively handling such a workflow. As indicated earlier, it was decided to use the Sharepoint platform offered in the cloud version of MS Office available to all University employees, but also to a certain extent the Alma library system and folder structures opened for this purpose on a dedicated server NAS (Network Attached Storage), as well as a paper register to control the transfer of original documents.
To cope with these challenges, it was decided to adopt an iterative solution, allowing for gradual solving of problems and getting closer to a production version that is understandable and acceptable to all stakeholders. The key points were two pilots carried out in 2023, but due to the complexity of the task, it was assumed that the first period after the implementation of the entire process would also be treated as the next stage of development and refinement of the workflow.
@@
In practice, it turned out that in the first period after implementation, a bundle of three parallel paths of different lengths was de facto formed, connecting with each other at specific points (which is presented in a simplified and synthetic way in Fig. 7). Path “A” refers to the handling of original manuscripts, the transfer of which is controlled analogously using paper documentation containing signatures confirming the transfer and receipt of objects by the people involved. Path “B” is the main axis of the entire project – it refers to the “life” of Excel files and intended for collecting and storing (at this stage of the project) metadata describing individual manuscripts. This path is implemented entirely in the digital environment within the structure of folders created on a dedicated Sharepoint website. The bundle is completed by path “C”, which includes activities related to the scanning process and is closely related to both paths mentioned above.
@@
It was assumed that the entire process would be continuous and systematic, and that the workload at its individual points would be adjusted so that the flow would be smooth, and the workload of individual people involved in the process would be more or less constant. However, it turned out that in practice a different solution was developed, perhaps more effective from the point of view of using the resources of the Jagiellonian Library. Currently, work is performed partly in a “batch mode”. The flow is paused at the place marked in Fig. 7 as “Creating batches”, a larger number of finally accepted files with metadata are collected and only such a “batch” is processed further. It is then processed as one action, after which work in this part of the workflow is suspended and waiting for the next “batch”. At the same time, in the first part of the workflow, which refers to the creation of metadata, work progresses continuously.
At this stage, it is necessary to perform analyzes and make decisions regarding the direction of further development of the workflow. Should the resources used be optimized in such a way as to achieve the previously assumed continuity of work within the entire workflow? Or should we adopt a hybrid solution in which metadata creation is carried out continuously and the scanning “factory” operates in a cyclical mode based on periodically processed larger “batches” and the second part of the workflow, work organization and resources involved should be adapted to such a solution? However, such a decision must be made considering the broader context of the Jagiellonian Library’s operation and its current and future involvement in other digitization projects.
7 Conclusions and future work
@@
When joining as a group of knowledge engineers and computer scientists in cultural heritage initiatives at the Jagiellonian University, we hit upon the challenge of preparing tools to collect metadata in a short period of time, with very limited human and financial resources and for a group of domain experts, some of whom are digitally excluded. It was important to ensure the scalability of the solution to hundreds of thousands of manuscripts and to take care of the validation of the metadata to be able to create a functional knowledge graph-based library out of it in the future. To achieve this, we designed a solution based on Microsoft 365 cloud, using MS Excel files as data entry interfaces, Office Script scripts for data validation and a dedicated folder structure in MS Sharepoint for file storage, and then together with the Jagiellonian Library, we formulated a workflow that combines existing systems and procedures with the proposed tools. The workflow has already been evaluated in two pilots in the DiHeLib project focused on digitizing the so-called “Berlin Collection” and in two workshops with international guests, which allowed for its refinement and confirmation of its correctness and usability for JL.
As future work, we plan to further evaluate the proposed workflow in other cases, including the Cultural Heritage Exploration and Retrieval with Intelligent Systems at Jagiellonian University (CHExRISH) project, where we will be able to examine the usability of the proposed solution at the JU museum, a different kind of institution in the GLAM group. We also plan to develop a knowledge graph-based catalog and related scripts to import metadata stored in Excel files, which is the ultimate goal of the reported work.
[…]
Acknowledgements
@@
[…]
Funding
This publication was funded by a flagship project “CHExRISH: Cultural Heritage Exploration and Retrieval with Intelligent Systems at Jagiellonian University” under the Strategic Programme Excellence Initiative at Jagiellonian University. The research for this publication has been supported by a grant from the Priority Research Area DigiWorld under the Strategic Programme Excellence Initiative at Jagiellonian University.
Declarations
[…]
Conflict of interest
@@
No datasets were created during this study.
[…]

### v2

@@ 4 Objectives and requirements
Our ultimate goal is to create a knowledge base that adheres to FAIR principles (Wilkinson et al., 2016), that is linked to general knowledge bases—such as Wikidata (Erxleben et al., 2014), Geonames [^13] https://www.geonames.org/., and YAGO (Rebele et al., 2016)—as well as to other cultural heritage bases—including the pan-European Europeana project (Haslhofer and Isaac, 2011) and bases describing the specific collections, e.g., the Kalliope catalog [^14] https://kalliope-verbund.info/en/. of approximately 600,000 records from 950 institutions located in German-speaking countries—and that will enable and facilitate a wide range of digital humanities research, including tasks done or supported by artificial intelligence tools.
As highlighted in the introduction, the objective of the work reported in this paper was to develop tools for metadata acquisition by domain experts and to combine them with existing systems and tools (described in Sect. 2) into a unified digitization workflow for cultural heritage collections. From the very beginning, the subset of collections stored in the Jagiellonian Library was chosen as the playground for the work. Therefore, our research concerns practical challenges of massive digitization (hundreds of thousands of documents) of manuscripts of immense historical value for the purpose of a digital library.
@@
    Tight time frame for tool development – four months were allocated for this task to be aligned with current digitization projects at Jagiellonian Library,
@@
    Lack of human and financial resources to develop dedicated software,
@@ 5 Methods and tools
    A set of Python scripts performing more detailed validation of metadata entered into the TKM and SEC (see Sect. 5.2),
@@
    The Arcarium Ontology (TAO) – interoperable data model for storing and retrieving rich metadata of manuscript collections housed at JU (see Sect. 5.3),
  8. 8.
    A set of Pythons scripts for populating knowledge stored in TKM and SEC into TAO (see Sect. 5.3),
  9. 9.
@@ 5.1 Designing TKM, SEC, and MAP
The Excel-based form for entering document metadata is the key data structure. First, it was assumed that each unit [^15] A set of documents marked with one shelfmark; in collection used for pilot study, one shelfmark was associated with all documents related to individual person. would correspond to one file, in which one of the sheets would be a metric containing the description of a unit as a whole (see Fig. 1), and the other would be used to collect metadata of individual documents represented by subsequent rows (see Fig. 2). In turn, columns of the latter sheet would represent successive fields in the document description. Thus, The Knowledge Matrix emerged. It was decided that each document would be classified into one of nine categories, a decision that determines the assignment of a particular set of fields to it. The list is a slightly modified version of the scheme used for the Varnhagen Collection (Kita-Huber and Jaglarz, 2022) and comprises the following categories:
@@
While the structure of the metric is very strict, as it contains fields necessary for creating records in the Alma library catalog (after conversion to MARC 21), the metadata collected for each document will be stored exclusively in the ontology and ultimately in a dedicated system that will enable their viewing. Therefore, it is more flexible and can be adapted to the specific nature of the resources being described. In the pilot study described, the metadata structure was created from scratch by a team of domain experts responsible for describing the documents (the whole data model can be explored in the shared ontology, see the Data availability section). Some basic fields are ascribed to all document types, e.g., “what is it / document title”, while other are associated with particular categories, e.g., “sender” and “recipient” are assigned to correspondence (categories 2 and 7), “issuer” is assigned to official documents (categories 4 and 5), and “author” to creative works (category 3). The description fields were further divided into mandatory (as “document number”) and optional (as “date remarks”). Detailed content restrictions were imposed on some of the fields, e.g., “date” field could only accept one of the specified date formats. The most important from the validation point of view (see Sect. 5.2) is “document number” filled with a string in format $x.y$, where $x$ is a digit of encoding document category of 1 to 9, and $y$ is a sequential number of a document within a unit.
@@
During subsequent iterations, the need for separate acquisition of person- and place-related metadata appeared. The problem arose from the fact that numerous persons and geographical objects were referred to in the documents under different names, resulting from variant spellings across languages or from the fact that a person or institution could bear different names at different stages of their existence. This need led to the creation of the Standard Entries Catalog. The SEC consists of four Excel files. Two of them are the parts of the actual catalog, presenting information about persons (see Fig. 3) and locations, respectively. Only the SEC coordinator is able to enter new data or modify existing entries, while other experts have read-only access. The other two files serve as collectors of proposals from domain experts: they can write suggestions of new entries or extending or modifying the existing entries in both main SEC files. The TKM records link to specific SEC entries by providing the SEC ID in a dedicated column. The SEC records are linked to external databases by providing URLs (e.g., Geonames for geographic names).
Finally, when the TKM file is finished, the validation script (see Sect. 5.2) extends it automatically with a MAP sheet designed to store bindings between unit’s card numbers and their scans. A single page may sometimes require more than just two scans (one for its recto, another for its verso side). There may be one or more paste-in slips attached to the page that require separate scanning. A telling example of this problem is one of the pages from Alexander von Humboldt’s Kosmos manuscript whose scanning produced as many as 14 different files (see pp. 62-75 in von Humboldt, 1845–1858). Therefore, to enable linking particular documents with their scans, a system of mapping card numbers to scan files—done manually by librarians in a separate MAP sheet—and mapping units with their documents to specific folder in a folder tree where the files are stored was introduced.
@@ 5.2 Data validation scripts
Data entered through interfaces (Excel files) must be evaluated to ensure the quality of the metadata is as high as possible. Two programs that perform data validation have been developed for this purpose: 1) Online validator, which validates data on the fly, as they are entered by the researcher into the TKM sheet (via an Office Script), and 2) Offline validator, which do its work after the sheet has been filled out completely – in this case, a full analysis is possible, which is technically impossible to perform at the Excel online script level (via a set of Python scripts). In both cases, the syntax of the input data and its completeness are analyzed. Offline validator additionally performs limited semantic validation including, but not limited to: verification whether the Author’s life span is not greater than 110 years, whether the entered dates actually existed in the calendar, whether the entered SEC IDs indeed exist in the SEC and correspond to the specified person/place. Feedback is provided to users through appropriate messages in a dedicated column A and cell background colors (in Online validator) and in the form of a report in a PDF file (in Offline validator).
Both programs were dedicated to data validation, however, during the course of work, their scope of operation was extended with additional functionalities resulting from the current needs:
@@
    Updating the TKM with new columns and fields, performed by Online validator (e.g., adding new fields that were not provided for earlier; the script performs this change without the need for additional user intervention);
@@
    Validation of SEC entries by Offline validator (an independent report is generated, visible to the SEC coordinator).
@@
For Offline validator, a dedicated virtual server on the JU infrastructure was set up and configured to perform a synchronization with Sharepoint once a day, run the validation scripts, and perform the synchronization again to propagate the validation reports.
5.3 TAO ontology population
The final step of the entire workflow is the ontology population, i.e., information about documents and individuals entered through the TKM and SEC interfaces is translated using custom scripts based on the owlready2 library into a knowledge graph in RDF notation and placed as instances in the TAO ontology. The result is a knowledge base built with rich metadata that allows querying in SPARQL to obtain answers to, e.g., (a) “To what shelfmark does a specific document belong?”, (b) “In what languages are the documents in a given shelfmark?”, (c) “Does a given person appear in the collection under other names, and if so, which ones?”. A detailed description of ontology design is presented in (do Valle Miranda et al., 2024).
6 Evaluation
@@
Data entry and validation are only the initial steps in an overall metadata acquisition process for digitization integrated with JL’s procedures. Working closely with JL, through a series of iterations, existing processes at JL were first identified, and then a new complete process was developed using Excel input interfaces, the Microsoft 365 cloud, and JL’s existing infrastructure. Importantly, the process was subjected to practical evaluation through two pilots and two workshops with international guests, which allowed for its improvement, and confirmation of its correctness and usability for JL.
@@
Designing, organizing, and implementing the workflow process, and consequently its implementation in practice, was a major organizational challenge due to three key factors. Firstly, it included the activities of three different groups of people, partly working in different places, with different levels of IT competences, using different IT tools on a daily basis and coming from different work and organizational cultures: (a) domain experts – humanities researchers, (b) computer scientists – researchers, academic teachers, doctoral students and practitioners, and (c) library workers. Secondly, the workflow was to be fed with different resources at different times: on the one hand, original manuscripts that require special protection and security procedures, and on the other hand, two types of digital documents: Excel files filled with metadata and scans of the manuscripts. Thirdly, the Jagiellonian Library did not have a ready-made tool capable of comprehensively handling such a workflow. As indicated earlier, it was decided to use the Sharepoint platform offered in the cloud version of MS Office available to all University employees, but also to a certain extent the Alma library system and folder structures opened for this purpose on a dedicated server NAS (Network Attached Storage), as well as a paper register to control the transfer of original documents.
@@
In practice, a bundle of three parallel paths was de facto formed, connecting with each other at specific points (which is presented in a simplified and synthetic way in Fig. 4). Path “A” refers to the handling of original manuscripts, the transfer of which is controlled analogously using paper documentation containing signatures confirming the transfer and receipt of objects by the people involved. Path “B” is the main axis of the entire project – it refers to the “life” of Excel files and intended for collecting and storing (at this stage of the project) metadata describing individual manuscripts. This path is implemented entirely in the digital environment within the structure of folders created on a dedicated Sharepoint website. Path “C” includes activities related to the scanning process and is closely related to both paths mentioned above.
@@
A practical assessment of the workflow has led to work being performed partly in a “batch mode”. The flow is paused at the place marked in Fig. 4 as “Creating batches”, a larger number of finally accepted files with metadata are collected and only such a “batch” is processed further. It is then processed as one action, after which work in this part of the workflow is suspended and waiting for the next “batch”. At the same time, in the first part of the workflow, which refers to the creation of metadata, work progresses continuously.
The bundle is completed by the execution of scripts that, based on archived metadata files, populate instances in the TAO ontology. During the pilot studies, this step was run manually, but ultimately it is planned to have routines that will call this step regularly to automatically populate the knowledge base with newly acquired metadata.
In the two pilot studies, the researchers created 235 TKMs and a number of records of individuals and places in the SEC. As the pilot studies were evaluating the continuous work of collecting metadata, according to the above workflow, their termination at a particular point in time meant that many TKMs were left in a draft version, which may contain errors or omissions. Therefore, during the ontology population step, it was assumed that any records with errors are skipped and errors are not analyzed, as they are mostly due to the specifics of the pilot studies (in the actual workflow, ontology population step will be done only on finished and validated TKMs).
As a result, a knowledge base consisting of 6084 axioms describing 852 individuals (people, documents, …) was created to evaluate the capabilities of the developed data model. The ontology has been made available in a public git repository available at https://gitlab.geist.re/pro/tao, along with a basic user manual to help familiarize oneself with the data we have collected. A detailed description of the data model can be found in (do Valle Miranda et al., 2024).
Evaluation of the entire workflow during two pilots and in two workshops with international guests confirmed the usefulness of the proposed solution. It allowed efficient collection of rich metadata by people with diverse backgrounds and integration with already existing digitization processes. The resulting knowledge base makes it possible to obtain answers to queries according to a pre-developed set of competency questions (cf. do Valle Miranda et al., 2024).
7 Conclusions
@@
When joining as a group of knowledge engineers and computer scientists in cultural heritage initiatives at the Jagiellonian University, we hit upon the challenge of preparing tools to collect metadata in a short period of time, with very limited human and financial resources and for a group of domain experts, some of whom are digitally excluded. It was important to keep in mind the scalability of the solution to hundreds of thousands of manuscripts and to take care of the validation of the metadata to be able to create a functional knowledge graph-based library out of it in the future. To achieve this, we designed a solution based on Microsoft 365 cloud, using MS Excel files as data entry interfaces, Office Script and Python scripts for data validation, a dedicated folder structure in MS Sharepoint for file storage, and a set of Python scripts for ontology population, and then together with the Jagiellonian Library, we formulated a workflow that combines existing systems and procedures with the proposed tools. The workflow has already been evaluated in two pilot studies and in two workshops with international guests, which allowed for its refinement and confirmation of its correctness and usability for JL.
Acknowledgement
@@
CRediT authorship contribution statement
Krzysztof Kutt: Writing – original draft, Writing – review & editing, Software, Methodology, Data curation, Conceptualization. Luiz do Valle Miranda: Writing – original draft, Writing – review & editing, Software, Validation. Jakub Gomułka: Writing – original draft, Methodology, Conceptualization. Grzegorz J. Nalepa: Writing – review & editing, Supervision, Resources, Project administration, Conceptualization, Funding acquisition.
Funding sources
This publication was funded by a flagship project “CHExRISH: Cultural Heritage Exploration and Retrieval with Intelligent Systems at Jagiellonian University” under the Strategic Programme Excellence Initiative at Jagiellonian University. The research has been supported by a grant from the Priority Research Area (DigiWorld) under the Strategic Programme Excellence Initiative at Jagiellonian University.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Writefully in order to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Declaration of competing interest
@@
The data model developed from the data collected in the pilot studies was made available in a publicly available git repository at: https://gitlab.geist.re/pro/tao.